# Universal Framework to Evaluate Automotive Perception Sensor Impact on Perception Functions

### v1

@@ Abstract
Current research on automotive perception systems predominantly focusses on either improving the sensors for data quality or enhancing the performance of perception functions in isolation. Although automotive perception sensors form a fundamental part of the perception system, value addition in sensor data quality in isolation is questionable. However, the end goal for most perception systems is the accuracy of high-level functions such as trajectory prediction of surrounding vehicles. High-level perception functions are increasingly based on deep learning (DL) models due to their improved performance and generalisability compared to traditional algorithms. Innately, DL models develop a performance bias on the comprehensiveness of the training data. Despite the vital need to evaluate the performance of DL-based perception functions under real-world conditions using onboard sensor inputs, there is a lack of frameworks to facilitate systematic evaluations.
This paper presents a versatile and cost-effective framework to evaluate the impact of perception sensor modalities and parameter settings on DL-based perception functions. Using a simulation environment, the framework facilitates sensor modality testing and parameter tuning under different environmental conditions. Its effectiveness is demonstrated through a case study involving a state-of-the-art surround trajectory prediction model, highlighting performance differences across sensor modalities and recommending optimal parameter settings. The proposed framework offers valuable insights for designing the perception sensor suite, contributing to the development of robust perception systems for autonomous vehicles.
@@ I Introduction
Deciding the best approacheciding the optimal perception sensor suite for an assisted and automated driving (AAD) vehicle remains an important, however, highly challenging task for the vehicle manufacturers. This is due to the dependency of most AAD functions on the availability of complete inputs for error-free operation. Recent state-of-the-art (SOTA) algorithms for high-level perception functions such as trajectory prediction and manoeuvre classification are increasingly based on deep learning (DL) models [13, 15, 25]. Most such models are developed and trained using ground truth datasets with no missing data. DL models have an inherent bias on the scope of the training data used during the development. Therefore, the quality and completeness of the inputs to these models play a crucial role in their successful deployment, questioning the performance integrity when such models operate using incomplete input data from onboard perception sensors.
Although the outputs of high-level perception algorithms directly impact the safe operation of the AAD features, research to evaluate the onboard perception sensor configurations that optimise AAD functional performance is almost non-existent [4, 21, 26]. The likely reason is the absence of suitable data to facilitate such studies since similar data logged using different sensor modalities are not readily available in most public automotive datasets [16]. In addition, the parameter settings of the perception sensors remain fixed during the data collection of naturalistic automotive datasets.
The existing approaches to evaluate and optimise the perception sensor suite for an AAD vehicle are threefold [3, 17, 21, 28]
@@
    Install as many and as high-spec perception sensors using prior engineering experience
@@
    Install the sensors as specified by the AAD algorithm developers
@@
    Statistical approaches based on existing sensor implementations on autonomous vehicle platforms
Neither approach supports the vehicle designers in selecting an optimal and cost-effective sensor suite for the vehicle platforms since the above approaches are either dependent on tier 1 supplier recommendations or based on empirical knowledge of the design engineers.
To address this knowledge gap, a universal framework developed using a virtual simulation environment is proposed, which enables exploring the impact of a given perception sensor configuration insert on a DL-based perception function. The framework offers maximum flexibility by adopting a modular structure, whereby virtual scenarios can incorporate a multitude of options. These include changes to the environmental conditions relating to both weather changes and road geometries as graphically illustrated in Fig. 1. The perception sensor configuration can be regarded as technology-agnostic, however, for generating synthetic data to conduct case studies, different sensor models can be used.
To demonstrate the utility of the proposed framework, the impact of the radar and the camera sensors is evaluated on a DL-based trajectory prediction model and the optimal setting for a common sensor parameter is determined. Although the case study focuses on surrounding vehicle trajectory prediction, the proposed framework is extendable universally towards other high-level perception functions such as manoeuvre classification and situational awareness.
@@
  • Presents a universal framework to evaluate the performance of perception functions based on different perception sensor configurations and different environmental conditions.
  • Offers a data-driven approach to specify the optimal perception sensors for an AAD function at the early stages of vehicle development.
  • Demonstrates the use of the proposed framework via a case study, where it evaluates the performance impact of different sensor modalities and parameter settings on a SOTA trajectory prediction model.
In addition, the ‘synthetic data generation’ stage of the proposed framework can produce data to improve the diversity of training and validation datasets. Existing real-world datasets can be augmented with synthetic data to introduce challenging environmental conditions or to compensate for the missing data. Such extended datasets can facilitate retraining models to uncouple domain biases, resulting in robust overall performances.
@@ II Related Works
Vehicles equipped with AAD features tend to have several established perception sensor modalities onboard such as radar, LiDAR, and vision sensors [19, 27]. The number of sensors and sensor modalities on a vehicle is likely to increase in conjunction with the ascending levels of automation defined by the Society of Automotive Engineers (SAE)[14]. Research on optimising an AAD perception sensor suite with or without linking to the performance of the perception algorithms is scarce[17]. It involves considering different perception sensor technologies and sensor parameters such as the range and the field of view (FOV), alongside factors including the environmental conditions applicable to the relevant operational design domain (ODD).
@@
Existing studies on optimising a vehicle’s perception sensor suite mostly concentrate on exploring a single sensor modality as opposed to comparing different modalities [24]. Most studies refer to ‘Perception sensor optimisation’ as the optimal placement of the available sensors to ensure maximum coverage of a predefined area. These approaches also lack the cost-benefit analysis of the impact of the sensor data towards the target perception functions [2, 24].
@@
T. Meng et al. present a framework for automatic selection and arrangement of the perception sensors using a genetic algorithm to solve for the optimal design of the sensor suite. The approach involves several prerequisites including the need to define optional sensors and positioning locations based on experience and feasibility, identify a critical point set (CPS) for emphatic detection, and predefine known ‘constraints’ based on prior engineering experience of the AAD feature. The CPS is based only on known edge cases, limiting the safety optimisation of the solution. Although the prerequisites intend to simplify the complexity of the genetic algorithm and the number of iterations, they act as a limitation demanding prior knowledge. This limits their approach to low SAE level features such as automatic park assist, which are commonly established using imperative algorithms. The opaque nature of DL-based algorithms makes their method unsuitable. [22].
@@
In [11], the authors introduce a simulation-based evaluation methodology to optimise the perception sensor setup in AAD vehicles. They use two virtual driving scenarios to demonstrate the usefulness of the method by analysing the impact of two different radar setups on low-level perception functions: object detection and tracking. The case study is limited to radars and lacks comparison between different sensor modalities. Instead of using sensor models to capture data, they use a development tool for environment perception applications, Baselabs Create, to synthesise the sensor data from the ground truths, i.e. no real perception sensor measurements or virtual sensor models were used during the simulation of the driving scenarios. Also, the sensor parameters were fixed during the synthesis process. Hence, their framework lacks the option to evaluate the sensor suite based on sensor parameters. The evaluation is only based on sensor mounting positions. In a follow-up publication, the authors apply their methodology to explore sensor positioning, specifically by altering the mounting heights to maximise the surround view coverage. However, again the study is limited to radars [12].
@@
In [9], the authors evaluate the perception performance on lane recognition as a function of the sensor data quality. They use different combinations of imaging sensors and camera lenses on lane recognition. However, this study was limited to cameras, and the performance analysis was based on detecting a single static element in the driving environment, ‘lane recognition’.
An established generic technique to optimise the overall sensing capability of a vehicle is sensor fusion. Many sensor-fusion algorithms have been published to leverage the individual strengths of different sensor modalities whilst compensating for each other’s weaknesses to optimise AAD functional performance [6, 1, 23]. However, the objective of most of these algorithms is to offer enhanced coverage and availability of sensor data; therefore, they lack direct associations with the requirements of the deployed perception algorithms, often leading to sub-optimal sensor configurations.
@@
T. Ma et al. introduce a novel metric to evaluate the sensor configurations of automated vehicles using both the perception algorithm performance and the selection of the sensors. Their evaluation metric, ‘perception entropy’, is calculated using Bayesian possibility theory to estimate the uncertainty of the detected targets and is formulated as a Gaussian distribution, conditioned on the sensor configurations. They link the definition of the metric to a selected perception algorithm, hence a change to the algorithm results in a change to the relevant sensor measurements and the relation to the algorithm’s performance.
@@
Another avenue of related research is adaptive sensor weighting techniques based on algorithmic outputs [18, 20]. This is where dynamically varying levels of importance are assigned to different sensor inputs based on the performance of the perception algorithms. Although a promising approach, the technique is applied after making the design decisions for the sensor suite. It also poses challenges in terms of accurately estimating the online algorithmic performance and stabilising the rapid changes in weights assigned to different sensors.
@@
In our view, optimising the sensor suite requires selecting the appropriate sensor modalities and configuring the sensor parameters to realise the optimal performance of the onboard perception algorithms. The sensors currently available in the market offer a range of specifications, often increasing in price with broader operational ranges. Therefore, the ability to perform impact analyses between different sensor configurations and the algorithmic performance of the perception functions is required. The proposed framework is a promising way to achieve the same and reshape the sensor optimisation landscape.
@@ III Problem Formulation
The perception sensor optimisation framework is formulated as a modular structure that brings together the essential elements for investigating the correlations between the configuration of the perception sensors and the performance of a given AAD perception function.
This section details the proposed framework using a simplified illustration of the vehicle’s sensing and perception system, specific to the trajectory prediction of surrounding vehicles as shown in Fig. 2.
@@
The interval $\left(i,j\right)$ depends on the manufacturer’s specifications of a given sensor for a selected ODD of an AAD function.
IV Methodology – Proposed Simulation Framework
Fig. 1 illustrates the proposed framework which implements the modular structure formulated in section III consisting of four stages: (i) synthetic data generation, (ii) data pre-processing, (iii) data utilisation, and (iv) data analysis.
@@ IV-A Synthetic Data Generation
A virtual environment is used to simulate the required driving scenarios for data generation. Following an empirical study, IPG CarMaker (CM) is chosen as the simulation tool for the framework due to its widespread use in the industry [10].
@@
The proposed framework offers maximum flexibility in choosing over a multitude of variable settings under two avenues. As shown in Fig. 3, firstly, the user can select the type of perception sensor and set the sensor parameter configuration in line with a test strategy. Secondly, they can select different vehicle types, manoeuvres, and environmental conditions when creating the scenarios, allowing them to incorporate edge cases, as necessary.
Execution of the tests and data logging is automated using a CM test automation feature, Test Manager. The simulation data is saved in a results file in the ERG file format.
@@ IV-B Data Preprocessing
Customised data preprocessing scripts are created in MATLAB for each sensor modality due to the differences in the recordable quantities for different sensor modalities. The script identifies the surrounding vehicles within a predefined boundary for the target vehicle and derives the input features for each ’vehicle of interest’ for the trajectory prediction model. It encapsulates the instantaneous road angle at each time frame to enable converting the data between the ego-centric frame of reference and the target vehicle’s frame of reference and generates the input data files for the trajectory prediction model. The ground truth trajectory for the target car is obtained directly from the simulation tool.
@@ IV-C Data Utilisation
IV-D Data Analysis
The metric root mean squared error (RMSE) is used for analysing the data. The RMSE values calculated on the predicted trajectories from the model using both the ground truth data and the sensor data are plotted against the prediction time horizon to compare the impact of different sensor modalities and parameter settings.
@@ V Case Study: Surround vehicle Trajectory Prediction
The merits of the framework are demonstrated using a DL-based model for trajectory prediction of surrounding vehicles. Six driving scenarios are created with a target vehicle performing a lane change manoeuvre on a highway. Lane change is identified as one of the safety-critical manoeuvres causing a high number of road accidents on highways [7, 8]. Therefore, developing knowledge on the impact of perception sensor configuration on lane change prediction performance is of high value.
V-A Model Selection

### v2

# Evaluation Framework for Sensor Configuration Impact on Deep Learning-Based Perception

@@ Abstract
Current research on automotive perception systems predominantly focusses on either improving the performance of sensor technology or enhancing the perception functions in isolation. High-level perception functions are increasingly based on deep learning (DL) models due to their improved performance and generalisability compared to traditional algorithms. Despite the vital need to evaluate the performance of DL-based perception functions under real-world conditions using onboard sensor inputs, there is a lack of frameworks to implement such systematic evaluations.
This paper presents a versatile framework to evaluate the impact of perception sensor modalities and parameter settings on DL-based perception functions. Using a simulation environment, the framework facilitates sensor modality selection and parameter tuning under different operational design domain conditions. Its effectiveness is demonstrated through a case study involving a state-of-the-art surround trajectory prediction model, highlighting performance differences across the sensor modalities radar and camera. Different settings for the parameter, horizontal field of view (HFOV) were evaluated to identify the optimal configuration. The results indicate that a radar sensor with a narrow HFOV is the most suitable configuration for the evaluated perception algorithm. The proposed framework offers a holistic approach to the design of the perception sensor suite, significantly contributing to the development of robust perception systems for automated driving systems.
@@ I Introduction
Deciding the optimal perception sensor suite for an automated driving system (ADS) remains an important, however, highly challenging task for the vehicle manufacturers. This is due to the dependency of most ADS perception functions on the availability of complete inputs for error-free operation. Most state-of-the-art (SOTA) high-level perception functions such as manoeuvre classification and trajectory prediction are increasingly based on deep learning (DL) models [15, 18, 31]. These models are often developed using complete ground truth datasets. The inference performance of a DL model has an inherent bias towards the scope, quality and completeness of its development dataset [36, 33]. Therefore, the performance integrity of DL-based perception functions when operating with incomplete input data from onboard perception sensors remains a crucial gap to be explored for both safe and efficient operation of ADSs.
Although the outputs of high-level, DL-based perception algorithms directly impact the safe operation of the ADS functions, research to evaluate the impact of onboard perception sensor configurations on the functional performance is rare [5, 26, 32]. The likely reason is the absence of suitable data to facilitate such studies since similar data logged using different sensor modalities are not readily available in most public automotive datasets [19, 29]. In addition, the parameter configurations of the perception sensors remain fixed during the data collection of naturalistic automotive datasets.
Some common approaches to determine the perception sensor suite for automated vehicles include the following [4, 22, 26, 37].
@@
    Base the sensor selection on empirical knowledge and prior engineering experience.
@@
    Select the sensors as specified by the ADS feature developers and/ or Tier 1 suppliers.
@@
    Statistical approaches based on existing sensor implementations on autonomous vehicle platforms.
Such initial sensor suite recommendations are then subjected to comprehensive simulation studies focused on architectural feasibility, safety-critical scenario coverage, and compliance with regulatory and packaging constraints. However, they lack the inherent link to the performance impact on the perception functions. Hence, these approaches rarely support the original equipment manufactures (OEMs) to select the most effective and cost-efficient sensors for the optimal performance of the DL-based perception functions deployed in a vehicle.
To address this knowledge gap, a framework that facilitates exploring the impact of different perception sensor configurations on the performance of a given DL-based perception function is proposed. Developed using a virtual simulation environment, it offers maximum flexibility by adopting a modular structure, whereby virtual scenarios can incorporate a multitude of options. These include different sensor types, parameter settings, manoeuvers, changes to the environmental conditions such as weather changes and road geometries as illustrated in Fig. 1. In doing so, the framework allows users to create scenarios that represent different operational design domains (ODDs) relevant to the perception function under evaluation, ensuring that sensor recommendations are aligned with the applicable ODDs.
The framework assumes the sensor is fixed at its standard location in current automated vehicles although the simulation tool allows for sensor placement at different positions on the vehicle body[16]. The perception sensor configuration can be regarded as technology-agnostic, however, for generating synthetic data to conduct case studies, sensor models from currently available sensors are used.
Existing literature on sensor optimisation only focuses on selecting and positioning sensors to optimise basic perception tasks, such as detection and classification. However, the proposed framework targets high-level DL-based perception functions at or beyond SAE Level 4 [17], where proactive planning replaces reactive planning and control, used in driver assistance systems such as Cruise Control (ACC) and Lane Keep Assist (LKA).
For DL-based perception models, the impact of sensor data quality on model performance is non-deterministic, making the framework to adopt an experimental approach. Therefore, it does not aim to optimise the entire sensor suite for an automated vehicle at once but serves as a tool to iteratively identify the most suitable sensor configurations for the each deployed DL perception function. Such knowledge enables the OEMs to make informed decisions on the overall sensor suite based on the deployed perception functions.
To demonstrate the utility of the proposed framework, the impact of radar and camera sensors is evaluated on a DL-based trajectory prediction model and the optimal setting for a common sensor parameter is determined. Although the case study focuses on surrounding vehicle trajectory prediction, the proposed framework is applicable to other high-level perception functions such as manoeuvre classification and situational awareness.
@@
  • Presents a novel framework to evaluate the impact of different perception sensor configurations on the performance of DL-based perception functions.
  • Offers a data-driven approach to identify the optimal perception sensor modalities and parameter configurations for an ADS function at the early stages of vehicle development.
  • Demonstrates the use of the proposed framework via a case study, which evaluates the performance impact of different sensor modalities and parameter settings on a SOTA trajectory prediction model.
In addition, the ‘synthetic data generation’ stage of the framework can enhance the diversity of existing real-world datasets by augmenting them with synthetic data to introduce edge cases or to compensate for missing data. Such extended datasets can facilitate retraining models to uncouple domain biases, resulting in robust overall performances [36, 33].
@@ II Related Works
Automated vehicles tend to have several established perception sensor modalities onboard such as radar, LiDAR, and vision sensors [24, 34, 26, 16]. The number of sensors and sensor modalities on a vehicle seem to increase in conjunction with the ascending levels of automation defined by the Society of Automotive Engineers (SAE)[17, 26]. Research on optimising an ADS perception sensor suite with or without linking to the performance of the perception algorithms is scarce[22, 20]. It involves considering different perception sensor technologies and parameter settings such as range, resolution, and field of view (FOV), alongside factors including the environmental conditions applicable to the relevant ODD.
@@
Most of the existing studies on automotive perception sensor optimisation define the problem as the optimal placement of perception sensors to ensure maximum coverage of a predefined area/ volume. Additionally, they only concentrate on exploring a single sensor modality as opposed to comparing different modalities [30, 20, 3, 21]. These approaches also lack the cost-benefit analysis of how sensor data impact the performance of targeted perception functions [3, 30, 21].
A detailed survey by Kiraz et al. provides a structured overview on the selection and positioning of sensors based on two categories; onboard and off-vehicle, with the focus of achieving cooperative, connected, and automated mobility. It highlights the limited research on optimising onboard multi-sensor configurations [20].
@@
T. Meng et al. present a framework for automatic selection and arrangement of the perception sensors using a genetic algorithm to solve for the optimal sensor suite. The approach involves several prerequisites including the need to pre-list candidate sensor types and positions based on experience and feasibility, identify a critical point set for emphatic detection, and predefine known ‘constraints’ based on prior engineering experience of the ADS function. This limits their approach to low SAE level features, commonly established using imperative algorithms, such as automatic park assist. Hence the method is unsuitable for DL-based perception functions due to their opaque nature [27].
A framework called ‘VESPA’ to optimise the placement and orientation of multi-modal sensors is presented in [6]. It uses three design space exploration algorithms to generate a sensor configuration that subsequently gets evaluated from an associated group of performance metrics. However, the optimisation is limited to SAE Level 2 autonomy features.
@@
In [13], the authors introduce a simulation-based evaluation method to optimise the perception sensor positioning setup and demonstrate the same using two virtual driving scenarios to analyse the impact of radar setups on low-level perception functions. The use of a development tool, Baselabs Create to synthesise sensor data from the ground truths limits the method’s ability to evaluate a sensor based on varying parameter settings. In a follow-up publication, the authors apply their methodology to explore radar sensor positioning, specifically by altering the mounting heights to maximise the surround view coverage [14].
@@
In [11], the authors evaluate the perception performance on lane recognition as a function of the sensor data quality. However, the data quality is contingent upon various combinations of imaging sensors and camera lenses, and the study is constrained to evaluating the camera sensor’s capability in lane recognition. recognition’.
@@
In an interesting attempt to link the selection of sensors to the performance of the perception function, T. Ma et al. introduce a novel metric to evaluate the sensor configuration of an ADS. The evaluation metric, ‘perception entropy’, is calculated using Bayesian possibility theory to estimate the uncertainty of the detected targets and is formulated as a Gaussian distribution, conditioned on the sensor configurations. The definition of the metric is linked to a selected perception algorithm, hence a change to the algorithm results in a change to the relevant sensor measurements and the algorithm’s performance [22].
In general, an established technique to optimise the overall sensing capability of a vehicle is sensor fusion. Many sensor-fusion algorithms have been published to leverage the individual strengths of different sensor modalities whilst compensating for each other’s weaknesses to optimise an ADS functional performance [8, 1, 28]. However, the objective of most of these algorithms is to offer enhanced coverage and availability of sensor data; therefore, they lack direct associations with the requirements of the deployed perception algorithms, often leading to sub-optimal sensor configurations.
@@
Another related research avenue is adaptive sensor weighting techniques based on algorithmic outputs [23, 25]. This is where dynamically varying levels of importance are assigned to different sensor inputs based on the performance of the perception algorithms. Although a promising approach, the technique is applied after making the design decisions for the sensor suite. It also poses challenges in terms of accurately estimating the online algorithmic performance and stabilising the rapid changes in weights assignment during real-time operation.
@@
In our view, optimal sensor configuration covers a much broader scope than determining the positions and orientations of sensors, but selecting the appropriate sensor modalities and configuring their parameters to realise the optimal performance of the onboard perception functions. Due to the varying vehicle body shapes and the diverse options in DL-based perception algorithms for vehicle deployment, to achieve the optimal sensor suite, customised evaluations need to be performed to learn the interdependencies between algorithmic performance and sensor configurations. In addition, the perception sensors currently available in the market offer a range of specifications, often increasing in price with the increase in operational range. Therefore, the proposed framework, which offers an approach to evaluate the impact between different sensor configurations and the algorithmic performance of a perception function is a promising way to achieve value-adding sensor optimisation.
@@ III Problem Formulation
The evaluation framework is formulated as a modular structure that brings together the essential elements for investigating the correlations between the configuration of the perception sensors and the performance of a given ADS perception function. This section details the proposed framework using a simplified illustration of the vehicle’s sensing and perception system, specific to the function, trajectory prediction of surrounding vehicles as shown in Fig. 2.
@@
The interval $\left(i,j\right)$ depends on the manufacturer’s specifications of a given sensor for a selected ODD of an ADS function.
IV Methodology – Proposed Evaluation Framework
Fig. 1 illustrates the proposed framework which implements the modular structure formulated in Sec. III consisting of four stages: (i) synthetic data generation, (ii) data pre-processing, (iii) data feeding for inference, and (iv) performance evaluation
@@ IV-A Synthetic Data Generation
A virtual environment is used to simulate the required driving scenarios for data generation. The use of simulation enables reproducibility and consistency across experiments, which would be very difficult to achieve with real sensor data under real driving conditions. However, the framework is designed to be both modular and robust, to accommodate real-world data as it becomes available. Following an empirical study, IPG CarMaker (CM) is chosen as the simulation tool due to its widespread use in the industry [12].
@@
The evaluation framework provides maximum flexibility in data generation, allowing users to choose from a variety of variable settings across two avenues: (i) select the perception sensor modality and configure the parameters in line with a test strategy, (ii) select different vehicle types, manoeuvres, and environmental conditions in creating the scenarios, allowing to incorporate edge cases, as necessary. Thereby, it enables users to define scenarios that reflect various ODDs pertinent to the perception task being assessed.
Execution of the tests and data logging is automated using Test Manager, a test automation feature in CM. The logged data during simulation get saved in a results file in ERG file format.
@@ IV-B Data Preprocessing
Custom data preprocessing scripts are developed in MATLAB for each sensor modality, accounting for variations in the recordable data based on the different sensor types. The script identifies the surrounding vehicles within a predefined boundary for the target vehicle and derives the input features for each ‘vehicle of interest’ for the prediction model. It encapsulates the instantaneous road angle at each time frame to enable converting data between the ego-centric frame of reference and the target vehicle’s frame of reference, and generates the input data files for the trajectory prediction model. The ground truth trajectory for the target vehicle is obtained directly from the simulation tool.
@@ IV-C Feeding for Inference
IV-D Performance Evaluation
The metric to use for performance evaluation is dependant on the perception function being studied to determine the optimal sensor configuration. The metric/s used during the DL model development for training and evaluation are suitable candidates. The baseline metric value needs to be established to facilitate comparison between the function’s performance against parameter variations, which is based on the ground truths. Alternatively a case study specific customised metric can be defined inline with [22].
@@ V Case Study: Surround vehicle Trajectory Prediction
The merits of the framework are demonstrated using a DL-based model for trajectory prediction of surrounding vehicles. It is used to determine the appropriate sensor modality based on the HFOV and subsequently narrow down the optimal parameter value of HFOV for the function.
V-A Optimisation Scope and Assumptions
Sensor selection is a problem based on the optimisation of a multitude of factors including the accuracy, precision, measurement resolution, range, coverage, power consumption, latency, cost, environmental robustness, etc. [20]. However, the performance of DL-based perception models is heavily influenced by the data quality. Therefore, the optimisation basis of this study is on sensor parameters such as ‘range’ and ‘HFOV’ for coverage and ‘resolution’. Whilst acknowledging the inherent interdependencies of these sensor parameters to jointly influence the perception performance, the limitations in sensor models within simulation tools make it challenging to capture these intricate influences with full fidelity. Therefore, to simplify the experiments, a controlled experimental design is adopted to isolate the effect of a single parameter on model performance.
In addition, to isolate the impact of sensor parameter variation, sensor placement is fixed at common positions found in current automated vehicles for the two sensors, as shown in Fig.4 [26, 20]. Since the considered trajectory prediction models are DL-based, sensitivity of the model’s performance to data quality remains unknown. Therefore, the optimisation is based on an experimental approach.
The ODD of the trajectory prediction models in the study is highway driving, under clear daylight conditions. Therefore, the resultant sensor recommendations are applicable to similar driving environments. In addition, the framework is specifically developed to evaluate the sensor setting’s impact on the performance of high-level, DL-based perception functions. The inputs to the perception models at these levels are typically derived from object-level features such as position and velocity rather than raw sensor data. Therefore, parameters like sampling rate and resolution hardly influence the model outcome, hence not varied in this study. The outcome of the case study is only applicable to DL-based trajectory prediction deployed at or above SAE level 4.
V-B Model Selection